OpenStudy (anonymous):

Use the discriminant to determine the number and type of solutions for the following equation. 24x2 – 13 = 0 A. Zero real solutions B. One rational solution C. Two rational solutions D. Two irrational solutions

OpenStudy (anonymous):

@johnweldon1993 please show me how to do this

OpenStudy (mathmale):

It's easy to recognize that your 24x2 – 13 = 0 has the basic form\[ax^2+bx+c=0.\]What are the values of a, b and c in your 24x2 – 13 = 0

OpenStudy (mathmale):

What does "discriminant" mean in this setting? What is the value of \[b^2-4ac\] in this particular problem? What does that tell you about the number and the type of solutions to 24x2 – 13 = 0?

OpenStudy (anonymous):

it doesnt have a c

OpenStudy (johnweldon1993):

Sure...as mathmale has shown above... Your equation takes the form of \(\large ax^2 + bx + c = 0\) 24x^2 - 13 = 0 Notice how 'a' is the coefficient of the x^2 ....and b is the coefficient of the x ....and 'c' is just the constant... so it looks like your equation has a = 24 b = 0 c = -13 The discriminant is the part of the quadratic equation that looks like \[\large b^2 - 4ac\] if the result is > 0 you have 2 real root if the result is = 0 then you have 1 real root if the result is < 0 you have 2 complex roots so what do you get when you plug in your numbers and solve that?